OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 parallel perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (anonymous):

put both equations in y = mx + b form and the m position is the slope... y = 2x - 8 -- slope here is 2 3x + 6y = 8 6y = -3x + 8 y = -3/6x + 8/6 y = -1/2x + 4/3 -- slope here is -1/2 now compare the slopes...parallel lines will have the same slope. Perpendicular lines will have negative reciprocal slopes. To find the negative reciprocal, you just " flip " the slope and change the sign. For example, the negative reciprocal of 6 is -1/6. You see how I flipped the slope and changed the sign. so you have 1 slope of 2 and another slope of -1/2.....what do you think ?

OpenStudy (anonymous):